# Q2E: Query-to-Event Decomposition for Zero-Shot Multilingual Text-to-Video Retrieval

### v1

@@ Abstract
Recent approaches have shown impressive proficiency in extracting and leveraging parametric knowledge from Large-Language Models (LLMs) and Vision-Language Models (VLMs). In this work, we consider how we can improve the identification and retrieval of videos related to complex real-world events by automatically extracting latent parametric knowledge about those events. We present Q2E: a Query-to-Event decomposition method for zero-shot multilingual text-to-video retrieval, adaptable across datasets, domains, LLMs, or VLMs. Our approach demonstrates that we can enhance the understanding of otherwise overly simplified human queries by decomposing the query using the knowledge embedded in LLMs and VLMs. We additionally show how to apply our approach to both visual and speech-based inputs. To combine this varied multimodal knowledge, we adopt entropy-based fusion scoring for zero-shot fusion. Through evaluations on two diverse datasets and multiple retrieval metrics, we demonstrate that Q2E outperforms several state-of-the-art baselines. Our evaluation also shows that integrating audio information can significantly improve text-to-video retrieval. We have released code and data for future research. [^1] https://github.com/dipta007/Q2E
@@ 2 Introduction
Figure 1: In Q2E, we extract prequel, current, and sequel events, along with an audio transcript and video description to enrich the query context and enhance video retrieval. This information matches against both visual, textual, and speech-based descriptions (matching phrases are highlighted in the same color) enabling the retrieval of the correct videos while effectively filtering out a visually similar but non-relevant video.
Widely used video retrieval websites can give the impression that video retrieval is a solved problem: a user types in a query, and gets back a nicely curated list of relevant videos. However, such approaches can leverage rich metadata, or subsequent human-creator curation (such as titles, or search-optimized descriptions), that go beyond the information conveyed directly in that video. For example, as shown in Fig. 1, there are many different videos of fires, but if a user provides a basic query such as “2025 LA fire” in order to find videos providing information about that specific fire, how can a system satisfy this request? In this paper, we present an approach we call Q2E (query-to-event decomposition) that leverages the embedded knowledge that LLMs have about complex events, and use this knowledge to automatically decompose and elaborate a user’s query, while simultaneously using this knowledge to identify salient visual and audio-based features from the video.
Recent work (Liu et al., 2024b) has found that users typically input basic queries, expecting the system to capture “fine-grained semantic concepts” and understand complex relationships between videos and text. Our proposed approach decomposes queries into multiple fine-grained events, providing critical elaboration of the implicit information requests of queries. Additionally, most current text-to-video retrieval systems are trained and evaluated on widely used datasets, such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011), which feature generic, high-level queries like “a person is explaining something” (a random query from MSR-VTT). These generic queries restrict text-to-video retrieval systems’ learning capacity, preventing them from effectively retrieving specific videos based on precise user queries, i.e., “2025 LA Fire”. Finally, most text-to-video retrieval datasets focus on English language videos, whereas videos transcend linguistic boundaries, necessitating retrieval methods adaptable to other languages.
Motivated by these challenges, we explore how to use query expansion and video+audio features to retrieve highly specific videos across different languages. Furthermore, we explore how the extensive knowledge embedded in Large-Language Models (LLMs) and Vision-Language Models (VLMs) can be effectively transferred to text-to-video retrieval. To achieve this, we propose a novel zero-shot text-to-video retrieval method that integrates the pre-existing knowledge of LLMs and VLMs to enhance text-to-video retrieval.
@@
Underlying our approach are three key insights: (1) Transferring knowledge from LLMs through query decomposition enhances the understanding of coarse-grained human queries. This provides richer context and enables the retrieval of prequel and sequel videos that would otherwise be overlooked. (2) While VLM-based captioning and large speech ASR models can generally describe an image/video frame and transcribe a portion of speech, the outputs can be repetitive, too low-level, missing broader themes that occur across frames. However, we can use an LLM to remove redundancy, and automatically summarize these lower-level descriptions for empirical effectiveness. (3) In our approach, determining whether an individual video is relevant requires aggregating across multiple forms of similarity and relevancy judgments, each of varying strength. Entropy-based rank fusion provides an intuitive way of performing this aggregation and ranking, while surpassing both simple fusion methods (mean, max) and more complex approaches like Reciprocal Rank Fusion. Using these insights, Q2E achieves strong performance gains on two varied and challenging datasets.
In summary, our contributions are as follows:
@@
  • We propose a novel framework that leverages the embedded world knowledge of LLMs to enrich human queries with prequel, current, and sequel-based events.
@@
  • Our zero-shot method improved performance on the widely used text-to-video retrieval dataset, MSR-VTT (Xu et al., 2016) (+4 NDCG points above our baseline), as well as on the more enriched, multi-lingual event-focused retrieval dataset, MultiVENT (Sanders et al., 2023) (+8 NDCG points).

### v2

@@ Abstract
Recent approaches have shown impressive proficiency in extracting and leveraging parametric knowledge from Large-Language Models (LLMs) and Vision-Language Models (VLMs). In this work, we consider how we can improve the retrieval of videos related to complex real-world events by automatically extracting latent parametric knowledge about those events. We present Q2E: a Query-to-Event decomposition method for zero-shot multilingual text-to-video retrieval, adaptable across datasets, domains, LLMs, or VLMs. Our approach demonstrates that we can enhance the understanding of otherwise overly simplified human queries by decomposing the query using the knowledge embedded in LLMs and VLMs. We additionally show how to apply our approach to both visual and speech-based inputs. To combine this varied multimodal knowledge, we adopt entropy-based fusion scoring for zero-shot fusion. Q2E outperforms the previous SOTA on the MultiVENT dataset by 8 NDCG points, while improving on MSR-VTT and MSVD by 4 and 3 points, respectively, outperforming several existing retrieval methods, including many fine-tuned and SOTA zero-shot approaches. We have released both code and data. [^1] https://dipta007.github.io/Q2E/
@@ 1 Introduction
Figure 1: In Q2E, we extract prequel, current, and sequel events, along with an audio transcript and video description to enrich the query and video context, respectively. This decomposed queries are matched across visual, textual, and speech-based descriptions (matching phrases are highlighted in the same color) enabling the retrieval of the correct videos while effectively filtering out a visually similar but non-relevant video.
Making sense of complex real-world events, such as a natural disaster or fire, requires understanding more than just one individual event or action. In addition to the multiple events that occur during that broader event (such as structures burning or evacuation), one might need to understand the events that lead up to it (such as dry conditions), or that arise from it (such as recovery and rebuilding). Videos present a rich medium for conveying this multi-faceted information, as they contain much more than what is visible at any given point in time: they can show changes and dynamic actions, and through an audio track, allow for natural language descriptions, either in-the-moment or after-the-fact, about that event. Being able to automatically retrieve videos about a queried event could provide the information necessary for users to better make sense of events they need to know about. However, there are multiple challenges to this: first, while a query might be a concise reference to a particular event, such as “2025 LA fire,” a user would still want all this information. Second, this information might be scattered across multiple videos, with any given video only showing a portion of the information needed. Third, these videos might not be in the language that the user knows.
We present a novel zero-shot method, Q2E, that addresses these problems. It consists of a proposed event decomposition method that leverages the embedded knowledge that LLMs have about those types of complex events and uses this knowledge to automatically decompose and elaborate (or expand) a query. We show how to use this knowledge to identify and score salient visual and multi-lingual audio-based features from the video, and combine this information to return the sought after videos.
@@
To achieve this, Q2E integrates the pre-existing knowledge of LLMs and VLMs to enhance text-to-video retrieval. Underlying our approach are three key insights: (1) Transferring knowledge from LLMs through query decomposition enhances the understanding of coarse-grained human queries. This provides richer context and enables the retrieval of videos showing what can lead to or result from the target event that would otherwise be overlooked. (2) While VLM-based captioning and large ASR models can generally describe an image/video frame and transcribe a portion of speech, the outputs can be repetitive, noisy, missing broader themes that occur across frames. However, a LLM based refiner can help to remove those noises and redundancies for empirical effectiveness. (3) In our approach, determining whether an individual video is relevant requires aggregating across multiple forms of similarity and relevancy judgments, each of varying strength. Entropy-based rank fusion provides an intuitive way of performing this aggregation and ranking, while surpassing both simple fusion methods (mean, max) and more complex approaches like Reciprocal Rank Fusion. Using these insights, Q2E achieves strong performance gains on three varied and challenging datasets (Chen and Dolan, 2011a; Xu et al., 2016; Sanders et al., 2023).
But isn’t video retrieval solved? Widely used video retrieval websites can give the impression that video retrieval is a solved problem: a user types in a query, and gets back a nicely curated list of relevant videos. However, such approaches can leverage rich platform-specific metadata or human-annotated data (like titles or search-optimized descriptions) that go beyond the information conveyed directly in that video. For example, as shown in Fig.˜1, there are many different videos of fires, but if a user provides a basic query such as “2025 LA fire” in order to find videos providing information about that specific fire, how can a system satisfy this request? Most current text-to-video retrieval systems are trained and evaluated on classic datasets, such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011a), which feature generic, high-level queries like “a person is explaining something” (a random query from MSR-VTT), rather than referencing a complex real-world event. Recent work (Liu et al., 2024b) has also found that users typically input basic queries, expecting the system to capture “fine-grained semantic concepts” and understand complex relationships between videos and text. Finally, Sanders et al. (2023) demonstrate the disparity that exists when retrieving non-English videos. Our work addresses these challenges by showing how to automatically expand the query with information needed to understand the event, and how to get video+audio features reflective of this information to retrieve highly specific videos across different languages. In doing so, we show how the extensive knowledge embedded in Large-Language Models (LLMs) and Vision-Language Models (VLMs) can be effectively transferred to text-to-video retrieval.
In summary, our contributions are:
@@
  • We propose a novel method that leverages the embedded world knowledge of LLMs to enrich human queries with prequel, current, and sequel-based events.
@@
  • Our zero-shot method improved performance on the widely used text-to-video retrieval dataset, MSR-VTT (+4 NDCG) & MSVD (+3 NDCG) as well as on the multi-lingual event-focused retrieval dataset, MultiVENT (+8 NDCG).
2 Q2E: Query-to-Event Decomposition
Figure 2: Q2E, the complete framework of our text-to-video retrieval system. The blue box represents the event decomposition module, while the green box illustrates the video decomposition module. The orange box represents our multi-layer Audio Decomposition module. The purple box fuses the ranks calculated from query-video, query-descriptions, and event-descriptions (event = prequel + current + sequel, description = multimodal description). In the non-ASR variant, the components within the orange box and orange dotted lines are excluded.
Modern dense retrieval approaches commonly compute an embedding-based similarity (e.g., cosine) between a query and each candidate; those scores rank candidates, and the top-scoring candidates are returned. For example, in video retrieval, one could embed the user’s query using an appropriate LLM, embed the video using an appropriate video embedder, and then compute the cosine similarity between those embeddings Sanders et al. (2023). This “query vs. video” similarity score is used to determine how relevant to video is to the query.
Q2E (shown in Fig.˜2) builds upon this general embedding similarity computation. To better capture the diverse and nuanced multimodal information conveyed in videos and human queries—including the event decompositions—we must extend those similarity computations. In order to capture the rich event information we get from our decomposition (§˜2.1), a single embedding of the video is not sufficient. We need to additionally get multimodal descriptions of the video (§˜2.2 and 2.3). Using these components, Q2E computes five types of scores $S_{i,v}$ for each video $v$: (a) a query vs. video score, as described above; (b) prequel, (c) current (d) sequel vs. multimodal-description score and (e) a query vs. multimodal-description score. While (a) helps to identify coarse-grained features such as “fire”, (b-e) provide fine-grained contextual details. These scores complement one another, which we merge through entropy-based fusion and ranking (§˜2.4).
2.1 Event Decomposition
Like any traditional text-video retrieval system, a user provides a natural language query, which can be a couple of words to 10+ words. In Q2E, we have used an LLM to decompose it into three sub-events: (1) Prequel, (2) Current, and (3) Sequel. Based on initial experiments, and to manage computation, we only considered five of prequel, current, and sequel events. We provide our template prompts for each in §˜A.6.
Prequel
“Prequel” refers to events that can lead to the current event described in the query. An event can have multiple prequels; for example, a wildfire can have strong wind, high-temperature alert, and dry lightning as prequels. Prequel events may enable the event in the query, but do not necessarily cause it.
Current
Sequel
“Sequel” events can be the results from the event described in the query. Like the prequel, one event can lead to multiple sequel events.
Decomposition Refinement
We refine these event decompositions with temporal, spatial, and primary event data. For this, we have used the same LLM to extract the temporal (2025), spatial (LA), and primary event (Fire) information. Although this rich information can be combined templatically, e.g., “P/S/C (2025, LA, Fire),” during initial development experiments, we found that refining it to read as a more natural, human-provided query markedly improved performance. We produced this refined query automatically using an LLM; we show this in §˜A.6.3.
2.2 Video Decomposition
In practice, we found that embeddings of videos / video frames were not sufficient to capture the varied actions described by our event decomposition through basic similarity computations. While alternative similarity approaches, such as cross-encoders, may be effective, they present computational challenges, requiring running the cross-encoder on every video/query pair. To address this, we instead extract captions from the videos themselves, with the aim that these descriptions would be sufficiently similar to our query event decomposition. We found two approaches to be effective for capturing complementary local and global elements across the video: (1) generating contextualized descriptions per selected frame, and (2) generating holistic video descriptions.
Contextualized Frame Descriptions
We uniformly sampled 16 frames from the videos (determined by trading off performance vs. computation— ablation in §˜A.3.3). We have tried uniform sampling and a scene-change-based sampling method  [^2] https://www.scenedetect.com/ for frame sampling, and found that uniform sampling, despite its simplicity, works better than scene-change-based sampling (ablation in §˜A.3.4). As the video contains temporal information and by design, video frames should be connected semantically, so rather than captioning each frame in isolation, we used a sliding window (window size=2) approach and provided the VLM model with the previous frame’s caption (when available). Conditioning on previous frame’s captions helped contextualize subsequent captions. Per video, we have K contextualized frame captions.
Video Descriptions
We found that current VLM models can perform dense captioning but miss the “bigger picture.” For instance, a VLM might describe a video of a wildfire with people running and fire everywhere but fail to recognize the higher-level conclusion of “wildfire.” To address this limitation, we provided an LLM model with all 16-frame descriptions from above and asked it to summarize them into a single video-dense caption, preserving the order of the frame captions. We carefully designed the prompt (§˜A.6.5) such that it preserves temporal information and focuses on broader picture rather than color, shape or objects.
2.3 Audio Decomposition
Speech is a critical component of many videos, especially when visual elements alone may not be sufficient enough to convey the full context. Despite using a multilingual ASR model (Radford et al., 2023), the multilingual videos posed a unique challenge to get consistent and accurate ASR transcriptions due to the background noise and different dialects of the real life videos in MultiVENT. To address this, we propose a multi-layer translation pipeline (Fig.˜2, left orange box). This pipeline leverages three models: (1) a multilingual ASR agent, whisper-v3 (Radford et al., 2023), which transcribes audio into both the original language and its English translation; 2) a translator agent, NLLB (Costa-jussà et al., 2022) to generate English transcriptions by translating the original transcript; and (3) a refiner agent, Llama-70B (Dubey et al., 2024), which refines both the English translations. Results showed all three are critical for sufficient transcription and translation quality (§˜A.3.2).
2.4 Scoring and Score Fusion
Query vs. Video
We utilized the original query and video to compute a similarity score. To do this, we follow Sanders et al. (2023) in adapting an image encoder: we uniformly sample and embed 16 frames, averaging their embeddings to generate the final video embedding. Cosine similarity was applied to compute the similarity score, which was then scaled to the 0–100 range to ensure comparability with other scores: $S_{i,v}|Q=100.0\times\operatorname{Sim_{v}}(Q,V),$ where $S_{i,v}|Q$ denotes the score vector for a given query $Q$, and $\operatorname{Sim_{v}}(x,y)$ returns the similarity score between query $x$ and video $y$ using an embedding model. Primarily we have used MultiCLIP (Delitzas et al., 2023) due to its superior performance with multilingual videos compared to the CLIP (Radford et al., 2021) model, although we show in our experiments that our approach successfully uses other image encoders too.
Model
Knowledge
Augmentation
[…]
MultiVENT
MultiCLIP - 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
MultiCLIP Event 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
MultiCLIP ASR + Event 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
InternVideo2-1B - 5.60 28.92 49.12 51.35 53.28 45.44 0.68 50.43 63.77 235.49 11
InternVideo2-1B Event 9.54 40.88 63.40 87.26 75.21 58.73 0.92 69.15 83.96 53.84 7
InternVideo2-1B ASR + Event 10.24 44.94 70.79 92.28 81.85 65.14 0.95 76.10 88.09 42.81 6
MSR-VTT-1kA
MultiCLIP - 43.52 69.05 76.88 43.62 13.85 7.71 0.54 59.72 54.27 20.29 2
MultiCLIP Event 44.52 71.26 79.40 44.62 14.29 7.96 0.56 61.51 55.84 17.91 2
MultiCLIP ASR + Event 46.23 73.37 81.71 46.33 14.71 8.19 0.58 63.59 57.83 18.73 2
InternVideo2-1B - 52.56 73.07 80.10 52.66 14.65 8.03 0.62 66.07 61.62 25.12 1
InternVideo2-1B Event 53.47 73.57 82.11 53.57 14.75 8.23 0.62 67.16 62.47 16.73 1
InternVideo2-1B ASR + Event 56.28 76.58 83.72 56.38 15.36 8.39 0.65 69.53 65.06 15.85 1
MSVD
MultiCLIP - 54.91 80.82 87.18 56.24 16.96 9.23 0.67 71.69 66.79 11.46 1
MultiCLIP Event 57.29 83.26 89.18 58.68 17.48 9.45 0.70 74.10 69.29 9.56 1
InternVideo2-1B - 63.59 84.85 89.46 65.03 17.85 9.49 0.74 77.51 73.68 12.32 1
InternVideo2-1B Event 63.77 85.30 89.99 65.24 17.94 9.54 0.74 77.84 73.95 9.40 1
Table 1: Performance comparison on MultiVENT, MSRVTT “1k-A” and MSVD dataset. Our method with Q2E, even without ASR, improves NDCG by 5–19 points on the MultiVENT dataset, 1-2 points on the MSRVTT dataset and 1-3 points on the MSVD dataset. Adding ASR further enhances performance, with an additional 3–7 points on MultiVENT and 2 points on MSRVTT, resulting in a total improvement of 8–26 points on MultiVENT and 3–4 points on MSRVTT, across two different encoders. Since the original MSVD videos are muted (Xu et al., 2016), we have excluded the ASR-based version. “Event” refers to the event decomposition (§˜2.1) and “ASR” refers to the audio decomposition (§˜2.3). Results from our whole method is highlighted.
Query vs. Multimodal Descriptions
In this approach, we used the original query but replaced the video content with the Multimodal Description Set, yielding a text-to-text similarity score instead of text-to-video. Specifically, we compute the maximum similarity between the query and all multimodal descriptions. Given $C$ descriptions associated with the video:
$S_{i,v}|Q=\max_{c=1:C}Sim_{t}(Q,Cap(v)_{c}),$ (1)
where $Cap(v)$ returns the multimodal description of the video $v$ and $Sim_{t}(x,y)$ computes the ColBERT [^3] We experimented with cosine-based SBERT (Reimers and Gurevych, 2019) and found that ColBERT outperforms SBERT in both accuracy and speed. Notably, the captions generated by LLMs may have some noise. ColBERT mitigates the noise’s impact by leveraging maximum aggregation at the token level, rather than compressing information into a single vector, which effectively reduces or ignores noisy elements in the multimodal descriptions. Additionally, prior studies (Warner et al., 2024) have shown that ColBERT’s approach minimizes information loss introduced by the averaging process used in SBERT, further enhancing its performance. similarity score between $x$ and $y$.
[…]
Arabic
     + Event 10.82 45.86 74.46 90.20 78.43 63.92 0.94 78.09 86.08 14.91 6.0
     + ASR 10.79 49.66 79.41 90.20 83.53 68.24 0.94 82.07 88.32 11.59 6.0
Chinese
     + Event 10.66 48.81 82.13 98.08 90.00 76.35 0.99 86.32 94.64 9.59 6.0
     + ASR 10.46 50.55 83.02 96.15 93.08 76.92 0.97 86.61 93.96 8.73 5.5
English
     + Event 10.61 51.66 86.45 100.00 97.69 82.50 1.00 90.57 98.04 8.07 5.0
     + ASR 10.61 51.66 87.80 100.00 97.69 83.85 1.00 91.43 98.12 7.74 5.0
Korean
     + Event 9.68 43.65 72.65 90.38 82.31 69.04 0.93 76.92 87.04 15.87 6.0
     + ASR 9.68 45.93 76.90 90.38 86.15 73.27 0.92 80.58 89.02 13.85 6.0
Russian
     + Event 10.39 50.55 81.61 92.31 89.62 73.46 0.96 84.62 91.46 10.88 6.0
     + ASR 11.21 53.49 85.40 98.08 94.23 76.54 0.99 88.70 95.02 9.13 5.0
Event Decomposition vs. Multimodal Descriptions
We use the refined prequel/current/sequel events from the query (§˜2.1) and the multimodal description set (§˜2.2) to compute a many-to-many similarity score. Specifically, we compute the maximum similarity score across all descriptions for each event and subsequently take the maximum similarity score among all the events. Formally:
$S_{i,v}|Q=\max_{l=1:L}\max_{c=1:C}Sim_{t}(E_{i}(Q)_{l},Cap(v)_{c}),$ (2)
This global maximum based approach reduces the negative impact of a single noisy decomposition by leveraging the global maximum instead of averaging, which is more vulnerable to noisy data. This global maximum effectively decrease the effect of plausible hallucination that can be expected with the LLM generation. For instance, for the query “LA Fire” one of the decomposed prequels might be “Man is playing with a dog” (a hallucinated event). As our method generates multiple decomposition of the same event but also takes the global maximum, Q2E effectively reduce the effect of hallucinated decompositions.
Score Fusion
A traditional approach of combining ranks from different experts is to train a fusion model. But as Q2E is based on zero-shot use of LLM agents, fusing all the scores presents a challenge. Naively averaging all scores would assume equal importance for each score, which does not reflect real-world scenarios. E.g., for videos relating to earthquakes, the Sequel vs. Descriptions score carries significantly more weight than the Prequel vs. Descriptions, as there is no visual indication of an earthquake before it occurs. Yin and Jiang (2024) demonstrated that entropy-based methods outperform others in zero-shot rank fusion tasks. In this entropy ranking, disparate similarity scores for an individual video are normalized to a softmax distribution, where low entropy indicates high confidence, while high inverse entropy corresponds to high confidence. Building on this insight, we use inverse entropy rank fusion to combine the scores.
$\hat{S}=\sum_{i}^{5}\frac{1}{H(P_{i})}*P_{i}.$ (3)
By applying Eq.˜3, we effectively assign greater weights to scores with lower entropy, emphasizing more confident predictions. This $\hat{S}$ contains an aggregate score for each video; $\hat{S}$ can be sorted to return the ranked list of videos for the query.
3 Results
Figure 3: Comparison of NDCG@10 scores across topic categories (Disasters, Political, Social, Technology) for five languages–Arabic, Chinese, English, Korean, and Russian. The results demonstrate the performance improvement from the baseline to our method (with audio or without), highlighting consistent gains across languages and domains.
3.1 Datasets and Implementation
We used the event-centric MultiVENT dataset (Sanders et al., 2023), the widely-used MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011b) dataset. We used the standard “MSR-VTT-1k-A” test split of 1,000 videos for computational constraint, as used in prior works (Jiang et al., 2022; Guan et al., 2023a; Jiang et al., 2023a).
Most of the MultiVENT videos are event-specific human queries with complex visual content (i.e., flood, people running) and rapid scene changes (i.e., news reports where multiple scenes are shown), where the average video is approximately 5.5 times longer than the MSR-VTT dataset (Table˜16). On the other hand, MSR-VTT and MSVD focus on general scenarios of everyday life, i.e., a man is playing with the dog. We have used different datasets to show the robustness of our method across different datasets of complexity. Full implementation details are in §˜A.5
3.2 Zero-Shot Text-to-Video Retrieval
We use MultiCLIP (per Sanders et al. (2023)) and InternVideo2 (the best-performing method on the MSR-VTT leaderboard [^4] https://paperswithcode.com/sota/zero-shot-video-retrieval-on-msr-vtt) as our baseline encoders. The comparison of our method with the baseline for MultiVENT, MSR-VTT-1kA and MSVD is shown on the Table˜1. We use the standard retrieval metrics: recall/precision at rank K, mean reciprocal rank (MRR), mean average precision (MAP), normalized discounted cumulative gain (NDCG), and mean and median rank (MnR, MdR).
Overall, combined with event decomposition and ASR, Q2E gives the best score in all the metrics across three different datasets and two different text-to-video encoders. But even without ASR, we outperformed the baseline in all metrics. To show the generalizability of our method across different text-video encoders, we applied it to both the MultiCLIP and InternVideo2 encoder.
Table˜1 shows the generality of our method. When applied to the InternVideo2, our method improved the MSR-VTT NDCG score by 1 point without audio and 4 points with audio. Additionally, the significant performance drop observed on the MultiVENT dataset underscores the dataset’s richness and complexity. Despite this drop, incorporating our method boosted the score by 19 points without audio and 26 points with audio, further validating its effectiveness.
Conversely, MultiCLIP achieved the best baseline score on the MultiVENT dataset (Sanders et al., 2023). Applying our method to MultiCLIP resulted in an 8-point NDCG improvement, setting a new state-of-the-art for the MultiVENT dataset. However, unlike InternVideo2, MultiCLIP does not use a cross-encoder reranker, which explains its comparatively lower performance on the MSR-VTT dataset. Our method still led to a 4-point improvement in the MSR-VTT NDCG score, demonstrating effectiveness across different encoders. Additionally, a full comparison with recent and popular baselines are provided in §˜A.1.
Next, we examine Q2E’s performance across languages (§˜3.3), the effect of different LLM sizes (§˜3.4), different rank fusion methods (§˜3.5) and various components of our method (§˜3.6). In the appendix, we examine the impact of VLM sizes (§˜A.3.1), different layers of ASR module (§˜A.3.2), number of frames (§˜A.3.3), and frame selection methods (§˜A.3.4). We provide a qualitative analysis in §˜A.2.
3.3 Performance Across Languages
MultiVENT includes five languages – Arabic, Chinese, English, Korean, and Russian. Table˜2 shows that building upon the MultiCLIP baseline, incorporating Q2E consistently improves retrieval performance across all metrics. outperforming the baseline by high numbers. We observed that audio-based decomposition plays a more significant role in languages like Arabic and Russian, whereas event-based decomposition is more crucial for Chinese and Korean. However, when combined, the two approaches complement each other, resulting in an aggregated performance boost.
To analyze the bias of the events and the effect of the bias on the performance of our method, we have used event category (provided in the main dataset) to divide the queries into 4 different buckets – Disaster, Political, Social, Technology. The results of different bucket across different language is provided in the Fig.˜3. We can see that our method improves the NDCG score irrespective of language or event category (except Arabic-Social). Also, the performance improvement is boosted with addition of audio features. The Social category, in particular, shows the most variation, with Arabic and Russian even experiencing a slight dip in performance, which could be linked to cultural differences in how queries are interpreted. These varying results across languages support the idea that query decomposition brings out language-specific nuances.
[…]
MultiCLIP 70.82 65.25 0.92 75.34
1B 78.71 72.28 0.95 82.50
3B 79.17 72.74 0.95 83.03
8B 79.04 72.59 0.95 82.91
70B (Q2E) 79.60 73.09 0.95 83.24
Table 3: Comparison of different LLM sizes within the Llama-3 family (1B to 70B parameters), specifically used for query decomposition and refining frame captions to video caption. Results are reported on the MultiVENT dataset. Full results in Table˜12, §˜A.3.5.
3.4 Effect of LLM size
Our original method used the Llama-3.3-70B LLM. While this model is highly robust and generalizable, its large size demands significant computational power and increased inference time. To demonstrate the robustness of our method, Table˜3 presents results for different LLM sizes, ranging from 1B to 70B parameters. Although the best performance is typically achieved with the 70B model, even the 1B model consistently outperforms the baseline NDCG by at least 8 points. These results demonstrate that our method can achieve comparable performance even when using significantly smaller LLMs, highlighting its practicality for real-world retrieval systems.
3.5 Effect of Rank Fusion Approaches
[…]
Neg. Exp. Ent. 67.23 61.97 0.93 73.20
RRF 70.91 65.29 0.93 76.29
Max 76.10 70.04 0.93 80.04
Mean 78.64 72.47 0.95 82.44
Inv. Ent. (Q2E) 79.60 73.09 0.95 83.24
Table 4: Comparison of Zero-Shot Rank Fusion methods: Neg. Exp. Ent. (Negative Exponential Entropy), RRF (Reciprocal Rank Fusion), Inv. Ent. (Inverse Entropy), Mean and Max. Results are reported on the MultiVENT dataset. Full results in Table˜13, §˜A.3.6.
Fusing different ranks in a zero-shot setting is a non-trivial task. Traditional approaches often involve training dense networks to combine scores (Yu et al., 2024) or leveraging LLMs to ensemble multiple expert outputs (Lu et al., 2024b, a; Jiang et al., 2023b). However, rank or score fusion is a significant challenge for LLMs, as they must process and compare a large volume of numerical data, including floating-point values (for scores) and discrete values (for ranks). In both cases, LLMs struggle to produce reasonable fusion ranks. We considered several aggregation methods, from simple techniques like Mean to more complex approaches such as Reciprocal Rank Fusion (RRF) and Inverse Entropy; due to space, we describe these formulations in §˜A.3.6. Results in Table˜4 indicate that even the simple Mean function performs well compared to RRF and Negative Exponential Entropy. Inverse Entropy demonstrated better performance, leading us to adopt it as the aggregation method.
While we adopted global max aggregation, for a fine-grained analysis of the max and mean operator, we have also tried different combination of max-mean aggregation over captions and events. The results are shown in the Table˜14. The results show that (1) Q2E performs the best compared with other aggregation methods. (2) Mean over captions underperformed the most, specifically learning nothing. We hypothesize that, as previously seen, the VLM models make noisy captions, and averaging over them hurts the model beyond what can be overcome. (3) Most of the other methods perform comparably with the global max, though we think global max wins because of its small performance gain and high simplicity. (4) The consistent performance across different evaluation metrics indicates that our method offers reliable improvements rather than metric-specific optimizations.
3.6 Effect of Different Components
[…]
Q2E 79.60 73.09 0.95 83.24
    w/o Video 67.74 62.43 0.93 73.96
    w/o Query 78.12 71.74 0.93 81.54
    w/o Event 77.77 71.47 0.94 81.75
Table 5: Analysis of different components in our method. Event = (Prequel + Current + Sequel). Results are reported on the MultiVENT dataset. The full result is reported on Table˜15 in §˜A.3.7.
Table˜5 presents an ablation study of event decomposition and multimodal description components of Q2E. As expected, the results show that the video is the most critical. Removing the query results in a 1-point performance drop, while removing events reduces performance by 2 points, highlighting the importance of events for retrieving relevant videos.